# Generative Point Cloud Registration

### v1

@@ Abstract
In this paper, we propose a novel 3D registration paradigm, Generative Point Cloud Registration, which bridges advanced 2D generative models with 3D matching tasks to enhance registration performance. Our key idea is to generate cross-view consistent image pairs that are well-aligned with the source and target point clouds, enabling geometry-color feature fusion to facilitate robust matching. To ensure high-quality matching, the generated image pair should feature both 2D-3D geometric consistency and cross-view texture consistency. To achieve this, we introduce Match-ControlNet, a matching-specific, controllable 2D generative model. Specifically, it leverages the depth-conditioned generation capability of ControlNet to produce images that are geometrically aligned with depth maps derived from point clouds, ensuring 2D-3D geometric consistency. Additionally, by incorporating a coupled conditional denoising scheme and coupled prompt guidance, Match-ControlNet further promotes cross-view feature interaction, guiding texture consistency generation. Our generative 3D registration paradigm is general and could be seamlessly integrated into various registration methods to enhance their performance. Extensive experiments on 3DMatch and ScanNet datasets verify the effectiveness of our approach. [Code]
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Paradigm comparison of our generative point cloud registration with conventional methods. Unlike geometry-only matching in previous methods, our approach introduces Match-ControlNet, a matching-specific 2D generative model that generates cross-view images pairs from point cloud data, providing rich color cues for enhanced geometric matching and pose estimation.
@@
Existing 3D registration methods can be roughly categorized into traditional approaches and data-driven deep methods. The traditional approaches include optimization-based fine alignment methods (besl1992method; yang2013go), which iteratively perform least-squares pose optimization for precise alignment, and handcrafted descriptor-based coarse alignment methods (rusu2008aligning; rusu2009fast), which capture local geometry to establish correspondences for hypothesize-and-verify registration. Deep registration methods, whether end-to-end (yew2020rpm; yew2022regtr) or descriptor-based (huang2021predator; qin2022geometric; jiang2023robust), exploit the power of deep neural networks to learn discriminative deep 3D features for robust matching. These deep methods significantly enhance the quality of estimated correspondences and improve registration accuracy.
@@
Despite the impressive performance achieved by current point cloud registration methods, their robustness remains limited in challenging scenarios that contain low overlap, repetitive patterns, or noisy points. Recent RGB-D registration studies (yuan2023pointmbf; mu2024colorpcr) have shown that incorporating rich texture and semantic cues from RGB images would significantly enhance the distinctiveness of point cloud descriptors, leading to improved matching accuracy. However, in geometry-only point cloud registration, the RGB images corresponding to the point clouds are unavailable, and existing methods rely solely on 3D geometric information for correspondence estimation and pose calculation. This raises an interesting question: “Can we still leverage color information to enhance geometry-only point descriptors for enhanced 3D registration?
@@
Motivated by this question and inspired by the recent successes of generative AI models (ho2020denoising; song2020score; yang2023diffusion; li2025controlnet; rombach2022high; Zhang_2023_ICCV; jiang2023se; wang2024freereg), we introduce Generative Point Cloud Registration, a new 3D matching paradigm that bridges the task gap between the 2D generative models and 3D matching tasks to enhance registration performance. Our key idea is to generate cross-view image pairs that are well-aligned with the corresponding source and target point clouds. These images provide rich color information to complement geometric features, enabling more robust matching (see Fig. 1). Unlike prevalent 2D generative models that focus on single image generation, our matching-specific image generation is pairwise. Importantly, to ensure high matching quality, the generated cross-view image pair should feature two key properties: 2D-3D geometric consistency and cross-view texture consistency. To achieve this, we introduce Match-ControlNet, a matching-specific, controllable 2D generative model. Match-ControlNet leverages ControlNet’s depth-conditioned generation capabilities to produce images geometrically aligned with depth maps (derived from the point cloud pairs), ensuring 2D-3D geometric consistency. Additionally, by incorporating coupled conditional denoising and coupled prompt guidance, Match-ControlNet enables effective cross-view image feature interaction, achieving mutual texture message passing and thereby enhancing cross-view texture consistency. Finally, we propose a zero-shot geometric-color fusion mechanism that leverages pretrained large vision models (e.g., DINOv2 (oquab2023dinov2) and Stable Diffusion (rombach2022high)) to extract discriminative zero-shot features of generated images for enhancing geometric descriptors via weighted concatenation.
It should be pointed out that our Generative Point Cloud Registration framework can operate in both zero-shot and few-shot settings (with minimal fine-tuning samples), each providing valuable color information to enhance precision. Moreover, our framework is general and can be integrated with various 3D registration methods to enhance their matching accuracy. Experiments on 3DMatch and ScanNet datasets validate the effectiveness of our proposed method. To summarize, our contributions are as follows:
@@
  • Unlike conventional single-image generation, we develop an effective Match-ControlNet for matching-specific, pairwise image generation. It incorporates depth-conditioned generation, coupled conditional denoising, and coupled prompt guidance to ensure that the generated image pairs maintain 2D-3D geometric consistency and cross-view texture consistency.
  • Our Generative Point Cloud Registration framework is general and plug-and-play. Benefiting from our effective zero-shot geometric-color feature fusion and XYZ-RGB fusion schemes, it can be integrated with various 3D registration approaches to provide free-lunch color information, enhancing their performance.
@@ 2 Related Work
Traditional 3D Registration Methods. Traditional point cloud registration methods are typically categorized into coarse and fine registration approaches. Iterative Closest Point (ICP) (besl1992method), a prominent fine registration method, iteratively computes nearest-neighbor correspondences and performs least-squares optimization for pose estimation. Go-ICP (yang2013go) enhances ICP’s robustness to initialization errors through a branch-and-bound (BnB) global search. Trimmed ICP (chetverikov2002trimmed) further improves robustness by optimizing over minimal subsets to handle outliers. Additional variants like (sharp2002icp; fitzgibbon2003robust; bae2008method; gressin2013towards; deng2018ppfnet) also demonstrate promising precision in fine alignment. Coarse registration methods generally combine handcrafted geometric descriptors with robust pose estimators, such as RANSAC. (johnson1999using) develops the spin image-based shape descriptors for surface matching and object recognition. USC (tombari2010unique) improves the feature descriptors using an shape context-aware unique local reference frame to improve matching accuracy. SHOT (salti2014shot) introduces a 3D histogram-based feature using normal vectors to describe surface. PFH (rusu2008aligning) and FPFH (rusu2009fast) constructs a discriminative and efficient local descriptor based on the oriented histogram with pairwise 3D representations. Other notable coarse methods, including (mohamad2014generalized; mohamad2015super; xu2019pairwise; huang2017v4pcs; ge2017automatic), have also achieved impressive registration precision.
@@
Learning-based Deep Registration Methods. Deep registration methods primarily consist of end-to-end approaches and deep descriptor-based methods. For end-to-end approaches, DCP (wang2019deep) introduces differentiable soft correspondences for SVD-based pose estimation. RPM-Net (yew2020rpm) incorporates the Sinkhorn layer and an annealing strategy to mitigate outlier inference. CEMNet and LatentCEM (jiang2021sampling; jiang2021planning) formulate the 3D registration task as a Markov decision process, and introduce a reinforcement learning-driven learning framework for planning-based trial-and-error pose searching (liu2024OPT; liu2023CIA; jiang2021action; jiang2022action). RegTR (yew2022regtr) designs an effective transformer-based correspondence regression module, addressing large-scale indoor scene registration in an end-to-end manner. For deep descriptor-based methods, 3DMatch (zeng20173dmatch) employs a Siamese 3D CNN to extract local geometric features for patchwise matching. FCGF (choy2019fully) develops a fully convolutional network to learn dense 3D features for pointwise matching. Predator (huang2021predator) introduces a cross-attention transformer between point cloud pairs for overlap perception and robust registration. GeoTransformer (qin2022geometric) integrates geometric embeddings into the transformer, enhancing feature discrimination. RoITr (yu2023rotation) designs a rotation-invariant transformation to further improve the rotational robustness of geometric descriptors. Other methods (wang2023zero; bai2020d3feat; li2022lepard; li2020iterative; choy2020deep; chen2023sira; fu2021robust; ao2023buffer) also demonstrate impressive performance in 3D registration. Beyond traditional and learning-based frameworks, this work introduces a new paradigm: Generative Point Cloud Registration. By integrating advanced 2D generative models with the 3D registration domain, our approach generates complementary color information for input point cloud pairs, producing color-enhanced geometric descriptors to improve precision.
Figure 2: Pipeline of Generative Point Cloud Registration. Given a source and a target point cloud, we first apply Match-ControlNet to generate their corresponding images. Next, we employ either zero-shot geometric-color feature fusion or XYZ-RGB fusion to create color-enhanced geometric descriptors, enabling high-quality correspondence estimation and robust pose estimation.
@@ 3 Approach
Problem Setting. Given a pair of source and target point clouds $\mathcal{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{3}\mid i=1,\dots,N\}$ and $\mathcal{Q}=\{\mathbf{q}_{i}\in\mathbb{R}^{3}\mid i=1,\dots,M\}$, point cloud registration seeks to recover their rigid transformation $\mathbf{T}=\{\mathbf{R},\mathbf{t}\}\in SE(3)$, comprising a rotation $\mathbf{R}\in SO(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$, to align them precisely. The optimal rigid transformation is typically computed by solving:
@@ 3.1 Motivation
Recent RGB-D point cloud registration methods (yuan2023pointmbf; mu2024colorpcr) have demonstrated that RGB images can significantly enhance geometry-only descriptors by providing rich color and semantic information. This enhancement facilitates the construction of higher-quality correspondences, leading to more robust registration.
However, in the context of 3D matching tasks, we focus on geometry-only registration using pure point clouds, as the relevant RGB data is unavailable. To overcome this limitation, we introduce Generative Point Cloud Registration, a general framework designed to generate high-quality RGB data for both point clouds, enabling geometric-color feature fusion for enhanced matching. Notably, unlike the conventional single-image generation focused by prevalent 2D generative models (rombach2022high; Zhang_2023_ICCV), our matching-specific image generation is pairwise, which should satisfy two key criteria:
@@
3.2 Zero-Shot Geometric Consistency Generation
In this section, we first address the 2D-3D geometric consistency generation through ControlNet (Zhang_2023_ICCV), a variant of Stable Diffusion (rombach2022high) with spatially localized image conditions (e.g. Canny edge maps and depth maps).
Stable Diffusion. Stable Diffusion is a widely used latent diffusion model for text-to-image generation. It operates within the latent space of a pretrained autoencoder, where a denoiser $\epsilon_{\theta}(\mathbf{x}_{t};t,\mathbf{c})$ (conditioned on the timstamp $t$ and tokenized text prompt $\mathbf{c}$) gradually refines the noisy latent feature $\mathbf{x}_{t}$ to clean one for image decoding. The denoiser follows a UNet architecture with an encoder, middle block, and skip-connected decoder, incorporating stacked transformer and residual modules. Each transformer module utilizes intra-image self-attention for contextual understanding and prompt-to-image cross-attention to guide generation.
ControlNet-driven 2D-3D Geometric Consistency. ControlNet further equips the denoiser of Stable Diffusion with a learnable encoder copy for encoding the conditional image $\mathbf{c}_{I}$, forming a conditional denoiser: $\tilde{\epsilon}_{\theta}(\mathbf{x}_{t};t,\mathbf{c},\mathbf{c}_{I})$. Consequently, the encoded condition features is concatenated with the original encodings of noisy latent representations $\mathbf{x}_{t}$ for conditional feature decoding via skip connections. Notably, ControlNet allows the use of depth maps as conditional inputs to generate RGB images that preserve geometric structures well-aligned with the provided depth prior. This capability perfectly aligns with our objective and motivates us to convert the source and target point clouds into their corresponding depth maps, $\mathbf{D}_{\mathcal{P}}$ and $\mathbf{D}_{\mathcal{Q}}\in\mathbb{R}^{H\times W\times 1}$, via the intrinsic matrix. Then, each produced depth map can be independently used to condition ControlNet to produce the geometrically consistent image.
Figure 3: Instead of independently performing ControlNet to generate source and target images, our Match-ControlNet integrates their denoising generation processes into a unified framework, facilitating feature interaction (i.e., mutual texture message passing) and enhancing their cross-view texture consistency.
3.3 Zero-Shot Texture Consistency Generation

### v2

# Geometry-to-Image Synthesis-Driven Generative Point Cloud Registration

@@ Abstract
In this paper, we propose a novel 3D registration paradigm, Generative Point Cloud Registration, which bridges advanced 2D generative models with 3D matching tasks to enhance registration performance. Our key idea is to generate cross-view consistent image pairs that are well-aligned with the source and target point clouds, enabling geometry-color feature fusion to facilitate robust matching. To ensure high-quality matching, the generated image pair should feature both 2D-3D geometric consistency and cross-view texture consistency. To this end, we introduce DepthMatch-ControlNet and LiDARMatch-ControlNet, two matching-specific, controllable 2D generative models. Specifically, for depth camera-based 3D registration with point clouds derived from the depth maps, DepthMatch-ControlNet leverages the depth-conditioned generation capabilities of ControlNet to synthesize perspective-view RGB images that are geometrically consistent with depth maps, ensuring accurate 2D–3D alignment. Additionally, by incorporating a coupled conditional denoising scheme and coupled prompt guidance, it further promotes cross-view feature interaction, guiding texture consistency generation. To address LiDAR-based 3D registration with point clouds captured by LiDAR sensors, LiDARMatch-ControlNet extends this framework by conditioning on paired equirectangular range maps projected from 360 LiDAR point clouds, generating corresponding panoramic RGB images. Our generative 3D registration paradigm is general and can be seamlessly integrated into a wide range of existing registration methods to improve their performance. Extensive experiments on the 3DMatch and ScanNet datasets (for depth-camera settings), as well as the Dur360BEV dataset (for LiDAR settings), demonstrate the effectiveness of our approach.
Index Terms:
Point cloud registration, large vision model, 2D generative model
@@ 1 Introduction
Figure 1: Paradigm comparison of our generative point cloud registration with conventional methods. Unlike geometry-only matching in previous methods, our approach introduces Match-ControlNet, a matching-specific 2D generative model that generates cross-view images pairs from point cloud data, providing rich color cues for enhanced geometric matching and pose estimation. In particular, we develop two variants: DepthMatch-ControlNet and LiDARMatch-ControlNet to handle depth camera-based and LiDAR-based point cloud registration tasks, respectively.
@@
Existing 3D registration methods can be roughly categorized into traditional approaches and data-driven deep methods. The traditional approaches include optimization-based fine alignment methods [4, 45], which iteratively perform least-squares pose optimization for precise alignment, and handcrafted descriptor-based coarse alignment methods [35, 34], which capture local geometry to establish correspondences for hypothesize-and-verify registration. Deep registration methods, whether end-to-end [48, 49] or descriptor-based [21, 32], exploit the power of deep neural networks to learn discriminative deep 3D features for robust matching. These deep methods significantly enhance the quality of estimated correspondences and improve registration accuracy.
@@
Despite the impressive performance achieved by current point cloud registration methods, their robustness remains limited in challenging scenarios that contain low overlap, repetitive patterns, or noisy points. Recent RGB-D registration studies [53, 30] have shown that incorporating rich texture and semantic cues from RGB images would significantly enhance the distinctiveness of point cloud descriptors, leading to improved matching accuracy. However, in geometry-only point cloud registration, the RGB images corresponding to the point clouds are unavailable, and existing methods rely solely on 3D geometric information for correspondence estimation and pose calculation. This raises an interesting question: “Can we still leverage color information to enhance geometry-only point descriptors for enhanced 3D registration?
@@
Motivated by this question and inspired by the recent successes of generative AI models [19, 38, 46, 24, 33, 55, 41], we introduce Generative Point Cloud Registration, a new 3D matching paradigm that bridges the task gap between the 2D generative models and 3D matching tasks to enhance registration performance. Our key idea is to generate cross-view image pairs that are well-aligned with the corresponding source and target point clouds. These images provide rich color information to complement geometric features, enabling more robust matching. Unlike prevalent 2D generative models that focus on single image generation, our matching-specific image generation is pairwise. Importantly, to ensure high matching quality, the generated cross-view image pair should feature two key properties: 2D-3D geometric consistency and cross-view texture consistency. To achieve this, we introduce DepthMatch-ControlNet and LiDARMatch-ControlNet, two matching-specific, controllable 2D generative models as shown in Fig. 1.
Specifically, DepthMatch-ControlNet is developed for the depth camera-based 3D registration, where the paired point clouds are derived from the depth maps. It leverages ControlNet’s depth-conditioned generation capabilities [55] to produce perspective-view RGB images geometrically aligned with depth maps, ensuring 2D-3D geometric consistency. Additionally, by incorporating coupled conditional denoising and coupled prompt guidance, it enables effective cross-view image feature interaction, achieving mutual texture message passing and thereby enhancing cross-view texture consistency. It should be pointed out that this version can operate in both zero-shot and few-shot settings (with minimal fine-tuning samples), each providing valuable color information to enhance precision. LiDARMatch-ControlNet is built for the LiDAR-based 3D registration that operates on the 360 LiDAR point clouds. It extends this framework by conditioning on paired equirectangular range maps projected from the LiDAR point clouds, and generates the corresponding panoramic RGB images with the geometric consistency and the cross-view texture consistency. Finally, we propose a zero-shot geometric-color fusion mechanism that leverages pretrained large vision models (e.g., DINOv2 [31] and Stable Diffusion [33]) to extract discriminative zero-shot representations of generated images for enhancing geometric descriptors via weighted concatenation. Moreover, our framework is general and can be integrated with various 3D registration methods to enhance their matching accuracy. Extensive experiments on the 3DMatch [54] and ScanNet [11] datasets (for depth-camera matching settings), and the Dur360BEV dataset [52] (for LiDAR matching settings) validate the effectiveness of our proposed method.
A preliminary version of this work was accepted by ICML’2025, and we expand upon it with comprehensive discussions and significant improvements. Specifically, our original manuscript focused solely on depth camera-based 3D registration using proposed DepthMatch-ControlNet. Beyond that, we now extend our Generative Point Cloud Registration framework to handle LiDAR-based 3D registration with point clouds captured by LiDAR sensors. To this end, we develop a novel and effective LiDARMatch-ControlNet. Here, we represent 360 LiDAR point clouds as equirectangular range maps, and innovatively take these range maps as condition to guide the generation of geometrically and texturally consistent panoramic images (see Sec. 3.4). To the best of our knowledge, this is the first successful exploration of LiDAR point cloud-to-panoramic image generation. Moreover, we provide a theoretical analysis of the coupled denoising diffusion mechanism used in both DepthMatch-ControlNet and LiDARMatch-ControlNet, showing that it effectively models the joint distribution of cross-view images and enables consistent generation (see Sec. 3.5). Finally, we supplement our work with extensive experimental comparisons on Dur360BEV dataset, verifying the effectiveness of our newly introduced designs (see Sec. 4.2).
To summarize, our main contributions are as follows:
@@
  • For depth camera-based 3D registration, we develop an effective DepthMatch-ControlNet for matching-specific, pairwise perspective-view image generation. It incorporates depth-conditioned generation, coupled conditional denoising, and coupled prompt guidance to ensure that the generated image pairs maintain 2D-3D geometric consistency and cross-view texture consistency.
  • For LiDAR-based 3D registration, we introduce a novel LiDARMatch-ControlNet, where we innovatively represent the 360 LiDAR point clouds as the equirectangular range maps to guide the generation of geometrically and texturally consistent panoramic RGB images. To the best of our knowledge, this is the first successful realization of LiDAR point cloud-to-panoramic image generation.
  • Our Generative Point Cloud Registration framework is general and plug-and-play. Benefiting from our effective zero-shot geometric-color feature fusion and XYZ-RGB fusion schemes, it can be integrated with various 3D registration approaches to provide free-lunch color information, enhancing their registration accuracy.
Figure 2: Pipeline of Generative Point Cloud Registration. Given a pair of source and target point clouds (captured by either a depth camera or a LiDAR sensor), we first employ the corresponding DepthMatch-ControlNet (LiDARMatch-ControlNet) to synthesize perspective-view (panoramic) RGB images. Next, we employ either zero-shot geometric-color feature fusion or XYZ-RGB fusion to create color-enhanced geometric descriptors, enabling high-quality correspondence estimation and robust pose estimation.
@@ 2 Related Work
Traditional 3D Registration Methods. Traditional point cloud registration methods are typically categorized into coarse and fine registration approaches. Iterative Closest Point (ICP) [4], a prominent fine registration method, iteratively computes nearest-neighbor correspondences and performs least-squares optimization for pose estimation. Go-ICP [45] enhances ICP’s robustness to initialization errors through a branch-and-bound (BnB) global search. Trimmed ICP [8] further improves robustness by optimizing over minimal subsets to handle outliers. Additional variants like [37, 14, 2, 18, 12] also demonstrate promising precision in fine alignment. Coarse registration methods generally combine handcrafted geometric descriptors with robust pose estimators, such as RANSAC. [22] develops the spin image-based shape descriptors for surface matching and object recognition. USC [40] improves the feature descriptors using an shape context-aware unique local reference frame to improve matching accuracy. SHOT [36] introduces a 3D histogram-based feature using normal vectors to describe surface. PFH [35] and FPFH [34] constructs a discriminative and efficient local descriptor based on the oriented histogram with pairwise 3D representations. Other notable coarse methods, including [29, 28, 44, 20, 16], have also achieved impressive registration precision.
@@
Learning-based Deep Registration Methods. Deep registration methods primarily consist of end-to-end approaches and deep descriptor-based methods. For end-to-end approaches, DCP [43] introduces differentiable soft correspondences for SVD-based pose estimation. RPM-Net [48] incorporates the Sinkhorn layer and an annealing strategy to mitigate outlier inference. RegTR [49] designs an effective transformer-based correspondence regression module, addressing large-scale indoor scene registration in an end-to-end manner. For deep descriptor-based methods, 3DMatch [54] employs a Siamese 3D CNN to extract local geometric features for patchwise matching. FCGF [10] develops a fully convolutional network to learn dense 3D features for pointwise matching. Predator [21] introduces a cross-attention transformer between point cloud pairs for overlap perception and robust registration. GeoTransformer [32] integrates geometric embeddings into the transformer, enhancing feature discrimination. RoITr [51] designs a rotation-invariant transformation to further improve the rotational robustness of geometric descriptors. Other methods [42, 3, 25, 23, 9, 6, 15, 1] also demonstrate impressive performance in 3D registration. Beyond traditional and learning-based frameworks, this work introduces a new paradigm: Generative Point Cloud Registration. By integrating advanced 2D generative models with the 3D registration domain, our approach generates complementary color information for input point cloud pairs, producing color-enhanced geometric descriptors to improve matching precision.
@@ 3 Approach
3.1 Problem Formulation
Given a pair of source and target point clouds $\mathcal{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{3}\mid i=1,\dots,N\}$ and $\mathcal{Q}=\{\mathbf{q}_{i}\in\mathbb{R}^{3}\mid i=1,\dots,M\}$, point cloud registration seeks to recover their rigid transformation $\mathbf{T}=\{\mathbf{R},\mathbf{t}\}\in SE(3)$, comprising a rotation $\mathbf{R}\in SO(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$, to align them precisely. The optimal rigid transformation is typically computed by solving:
@@ 3.1 Problem Formulation
Depth Camera-based Point Cloud Registration. In our original manuscript, we primarily target the depth camera-based 3D registration where the point clouds are derived from the depth maps as in [54, 11]. Due to the limited field-of-view, each point cloud typically captures only a partial observation of the scene. Thus, the resulting source and target point clouds (captured from different viewpoints) would suffer from partial overlap and incomplete geometry, increasing the challenge of accurate correspondence estimation and transformation recovery.
LiDAR-based Point Cloud Registration. Beyond the depth camera-based 3D registration describe above, our extended framework further generalizes the Generative Point Cloud Registration paradigm to LiDAR-based 3D registration with 360 point clouds captured by LiDAR sensors. In this setting, two temporally adjacent LiDAR sweeps are treated as the source and target point clouds for geometric matching. Due to the sparsity of LiDAR data, robust LiDAR matching remains a challenging problem.
@@ 3.2 Motivation
Recent RGB-D point cloud registration methods [53, 30] have demonstrated that RGB images can significantly enhance geometry-only descriptors by providing rich color and semantic information. This enhancement facilitates the construction of higher-quality correspondences, leading to more robust registration.
However, in the context of 3D matching tasks, we focus on geometry-only registration using pure point clouds, as the relevant RGB data is unavailable. To overcome this limitation, we introduce Generative Point Cloud Registration, a general framework designed to generate high-quality RGB data for both point clouds, enabling geometric-color feature fusion for enhanced matching. Notably, unlike the conventional single-image generation focused by prevalent 2D generative models [33, 55], our matching-specific image generation is pairwise, which should satisfy two key criteria:
@@
Given the significant disparities between depth camera-based point clouds and LiDAR-based point clouds (single-view vs. omnidirectional), as well as the distinct characteristics of their associated RGB images (perspective-view vs. panoramic), employing a unified point cloud-to-image generative model for both depth camera-based and LiDAR-based 3D registrations is inherently challenging. As such, we propose two variants: DepthMatch-ControlNet and LiDARMatch-ControlNet, designed to generate perspective-view RGB images for the depth-camera matching (see Sec. 3.3) and panoramic RGB images for the LiDAR matching (see Sec. 3.4), respectively.
3.3 DepthMatch-ControlNet
In this section, we first introduce the DepthMatch-ControlNet, which takes point cloud pairs derived from depth maps as input and generates the corresponding source and target perspective-view RGB images that exhibit strong 2D-3D geometric consistency and cross-view texture consistency for RGB-enhanced geometric matching.
3.3.1 Zero-Shot Geometric Consistency Generation
We first address the 2D-3D geometric consistency generation through ControlNet [55], a variant of Stable Diffusion [33] with spatially localized image conditions (e.g. Canny edge maps and depth maps).
Stable Diffusion. Stable Diffusion is a widely used latent diffusion model for text-to-image generation. It operates within the latent space of a pretrained autoencoder, where a denoiser $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t};t,\mathbf{c})$ (conditioned on the timestamp $t$ and tokenized text prompt $\mathbf{c}$) gradually refines the noisy latent feature $\mathbf{x}_{t}$ to clean one for image decoding. The denoiser follows a UNet architecture with an encoder, middle block, and skip-connected decoder, incorporating stacked transformer and residual modules. Each transformer module utilizes intra-image self-attention for contextual understanding and prompt-to-image cross-attention to guide generation.
ControlNet-driven 2D-3D Geometric Consistency. ControlNet further equips the denoiser of Stable Diffusion with a learnable encoder copy for encoding the conditional image $\mathbf{c}_{I}$, forming a conditional denoiser: $\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t};t,\mathbf{c},\mathbf{c}_{I})$. Consequently, the encoded condition features is concatenated with the original encodings of noisy latent representations $\mathbf{x}_{t}$ for conditional feature decoding via skip connections. Notably, ControlNet allows the use of depth maps as conditional inputs to generate RGB images that preserve geometric structures well-aligned with the provided depth prior. This capability perfectly aligns with our objective and motivates us to convert the source and target point clouds into their corresponding depth maps, $\mathbf{D}_{\mathcal{P}}$ and $\mathbf{D}_{\mathcal{Q}}\in\mathbb{R}^{H\times W\times 1}$, via the intrinsic matrix. Then, each produced depth map can be independently used to condition ControlNet to produce the geometrically consistent image.
Figure 3: Instead of independently performing ControlNet to generate source and target images, our DepthMatch-ControlNet integrates their denoising diffusion generation processes into a unified framework, facilitating feature interaction (i.e., mutual texture message passing) and enhancing their cross-view texture consistency.
3.3.2 Zero-Shot Texture Consistency Generation